# Distributed Multisensor ISAC

### v1

@@ V-C Recovery of the Correlation Reference Signal
Another approach to filter out remaining multipath is beamforming, which is discussed in the context of spatial precoding later on.
V-D Model based Bistatic Target Parameter Estimation
@@
This subsection deals with the estimation of target related parameters from observed ISAC OFDM frames. As described above, the 3D dynamic target state vector maybe estimated from a set of concurrent path parameters in delay and Doppler. This is not a restriction on the generality, as we can easily add beamforming and even directional estimation as will be discussed later. The aim here is to estimate the statistical delay Doppler distribution, which is also known as the 2D point spread function of linear time-variant systems. In Bello’s classical terminology [ref63] it is called Scattering Function (ScF). This is the magnitude squared of the Spreading Function (SF) $S(\tau,\alpha)$ that results from the inverse 2D Fourier transform of $H(f,t)$.
@@
$\chi(\tau,\alpha)=\int\limits_{-\infty}^{+\infty}x(t)x^{\ast}(t-\tau)\exp(\jmath 2\pi\alpha t)\mathrm{d}t$ (6)
The AF describes the resolution capability of the excitation waveform associated with its impulse compression performance and correlation gain. While the ideal AF according to (6) is a 2D Dirac delta function (”thumbtack”), for uniformly occupied OFDM frequency-time frames it appears as 2D $\operatorname*{sinc}$ function because of its limited bandwidth $B$ and frame duration $T$. Although the side lobes of this AF are often considered as disturbing, the main lobe is the narrowest and the correlation gain is the highest for a given area of support $BT$, in frequency-time domain. The resulting detector may be therefore seen as optimum in terms of maximum likelihood performance in case one path to be estimated, simple FFT processing, and constant noise energy distribution in delay domain.
@@
The first equality is the classical Wiener-Hopf equation. It is also known as Wiener-deconvolution. In the general framework of frequency domain system identification [ref30], estimation of the auto/cross power spectral densities $R_{X}(f)$ and $R_{Y,X}(f)$ of stationary random signals would need a separate statistical averaging procedure. However, it follows directly from the of OFDM paradigm that the orthogonality of the received OFDM carriers is kept and any leakage noise of the estimated samples in frequency domain is avoided. This results from the Cyclic Prefix (CPX), which is added before transmission and removed at the receiver before FFTs processing. The absence of further statistical averaging allows for rigorous simplification which finally leads to the well-known inverse filter approach which consist of dividing with the recovered correlation reference signal as summarized in equation (7) below. Actually, PCL applies a similar processing scheme [ref38]. This also corresponds to standard zero forcing equalization. Another reason that allows for factorization of frequency/delay and time/Doppler transform is related to the “underspread” characteristics of the mobile radio channel [ref95]. This means it is a slowly time-variant system, which can be handled as block wise time-invariant as long as the OFDM symbols are chosen to be short enough to avoid time-selective fading (the block fading assumption). This finally allows the factorization which avoids full 2D FFTs and matches perfectly to the OFDM receiver processing as depicted in Figure 15. The adjacent Doppler analysis is carried out by subsequent block of FFT along slow time over $M$ stacked CIRs. These FFT s can be also interpreted as a Doppler filter bank. The number of transformed consecutive OFDM symbols (the OFDM frame) determines the coherent radar integration time $T$. Further noncoherent integration can be added.
@@
Figure 15: Basic ISAC-OFDM receiver signal processing in frequency-time domain ans estimation of the scattering function.
Although the simple 2D FFT based frequency domain system identification scheme has several advantages, there are also some drawbacks and pitfalls. Among those are: (i) Even if the OFDM frame were completely filled with carriers of equal strength, the strong sinc sidelobes in the 2D delay/Doppler plane may mask weak reflections. The resolution of components of similar strength is limited by the AF mainlobe width which is given by the $1/B\times 1/T$ aperture size, where $B$ is the bandwidth and $T$ is the coherent processing interval in slow time. (ii) The estimated SF is calculated at the predefined FFT grid in the $(\tau,\alpha)$ plane. However, the maxima of $\left|S(\tau,\alpha)\right|^{2}$ which mark the $(\tau,\alpha)$ target parameters, are not bound to this grid. Therefore, we would need efficient local interpolation schemes if we apply the FFT. (iii) Dynamic targets that are revealed by a non-zero Doppler shift will also change their observed delay (resp. range). Therefore, the estimated SF may be be blurred by range Doppler migration if delay resolution (i.e. bandwidth) is high and the chosen coherent integration time is too long, since the Doppler filter bank assumes constant delays. (iv) The most fundamental drawback, however, comes into play when we are faced with a sparse, non-uniform distribution of the RE in the frequency/time frame that is so important for ISAC as discussed above. In this case, the AF is distorted in an uncontrolled manner, which may cause detection ambiguity and make correct estimation impossible. We should be aware, that application of non-rectangular window functions (tapering) in frequency and/or slow time aperture space are always suboptimal since that reduces resolution and increases noise induced variance. They do also not help at all in case of sparse RE distribution
In the sequel we describe an example that explains the problems of FT based estimation of the ScF that arise with aperture-limited and sparse observations in the frequency/slow time domain. At the same time, we introduce a model-based estimation approach, which can solve both problems. We consider a propagation situation which is determined by $P$ specular paths with delay/Doppler parameters $\tau_{p}$, $\alpha_{p}$ and path weight $\gamma_{p}$, respectively:
@@
where $H(f,t)$ denotes the slowly time-variant radio channel’s system response, limited to finite apertures, both in frequency and time. After a 2D Fourier Transform we obtain the complex SF $S(\tau,\alpha)$ which is affected by a sinc function in both dimensions. The symbol $\ast$ is a shortcut for a convolution operation in the dimensions $\tau$ and $\alpha$ respectively.
@@
Here, $\left|S(\tau,\alpha)\right|^{2}$ is the so-called ScF, where the detection takes place. It is also referred to as the statistical delay-Doppler distribution. For the example $P=2$, the red arrows in Figure 16 indicate ideal case of no bandwidth-time (BT) limitation in the $(f,t)$ aperture plane. But (8) and (9) also accounts for the the 2D rectangular limited support in the $(f,t)$ aperture domain that describes the observed data taken by the OFDM frame. If the OFDM carriers are uniformly power loaded this appears to be the optimal processing scheme. Some interpolation in delay-Doppler for better detection of maxima can be achieved by zero filling in the $(f,t)$ aperture space outside the observed OFDM frame. Although the two paths appear to be smeared in the delay-Doppler domain by the 2D convolution with a 2D $\operatorname*{sinc}$ function (see Figure 16 ), they can be easily identified since they are well separated.
@@
Now we consider a OFDM resource grid with sparsely allocated REs as indicted as an example by the red dots in the $(f,t)$-frame in Figure 18. The rippled surface is the magnitude of $H(f,t)$ according to (8) results from the superposition of the two equal power propagation paths. The complex valued $H(f,t)$ is effectively sampled by the sparsely allocated REs. The missing samples due to empty carriers are set to zero for FFT-processing. This is common practice since FFT requires a uniform sample grid. However, simply replacing samples that are not measured with zeros is a bad choice, as zeros are definitely wrong. The resulting estimated delay-Doppler distribution is therefore subject to corresponding distortions, which are characterized by uncontrolled side lobes and can make detection impossible as depicted by the example in Figure 17.
@@
Fortunately, the solution is obvious. A quick look at (8) explains that we can simply consider the equation given there for the analytical description of the problem as a parameterized model of the complex transfer function. So we just need to estimate the parameters of this model. If we have enough well allocated REs in the $(f,t)$ resource frame, we can estimate the model parameters of by fitting the $H(f,t)$ model at the measured RE positions. This is very well supported by the orthogonality of the system response in frequency domain with respect to the the OFDM carriers. Parameter estimation is achieved by minimizing a properly defined least squares cost function. In some general sense, this estimation procedure constitutes a multidimensional maximum likelihood parameter estimation scheme [ref64, ref67]. A deep learning based approach for joint delay-Doppler model parameter optimization from frequency and time samples can be found here [schieler2025prop_est_cnn]. Efficient real-time implementation is another problem and remains an active area of research. While batch processing may be not feasible, recursive estimation seems to be more appropriate. It is also worth mentioning that the model order $P$ is in most cases not known a priori and has to be determined. Finally, the complex path weights $\gamma_{p}$ is estimated. Moreover, as $\gamma_{p}$ is polarization sensitive, it can be given by a $2\times 2$ Jones matrix that accounts for HH, VV co-polar and HV, VH cross-polar I/O response.
@@
Model-based estimation allows us to obtain the set of discrete delay-Doppler parameters ($\tau$, $\alpha$) without having to apply an FFT. This avoids any aperture truncation and sparsity gap errors, which would reduce resolution and lead to spurious sidelobes, as discussed above. Figure 18 explains that the propagation data model effectively interpolates between the observed sparse samples in $(f,t)$, thus avoiding the artifacts that would result from a simple FFT application. At the same time, also the effect of high resolution is explained, since the model function implicitly extrapolates the observed data beyond the BT limits, where $H(f,t)$ is measured, thus evades the Fourier resolution limit (beyond Rayleigh resolution). A key aspect of this estimation procedure is that it decomposes the relevant propagation components in the complex linear signal domain. This way, their mutual interaction in the delay-Doppler domain because of limited support and sparsity in the observed frequency-time aperture domain is removed. The cancellation of overlapping non-resolved contributions and sidelobes in the complex signal domain seems to be more powerful than the well known CLEAN procedure [hoegbom1974clean] which works in the real-valued point spread or correlation domain.
@@
However, when considering the waveform of the data payload as an excitation signal for system identification, further considerations are necessary. Figure 19 shows samples of $H(f,t)$ at some cross section in $f$ or $t$. The problems are as follows: Empty REs have to be identified and excluded from the inverse filter operation. A null hypothesis test allows empty but noisy resource elements to be detected and removed from the estimation. Another problem relates to the non-uniform power level of data carriers in due to stronger CSI pilots and because of adaptive power loading and multi-level QAM coding. While the simple inverse filter would be affected by uneven noise levels, the SNR of the carriers can be estimated at the receiver and considered in parameter estimation as weighted least squares fitting which gives the samples with lower SNR a smaller weight.
@@
Figure 16: Observed delay-Doppler distribution. Frequency-time resource grid is limited in BT and fully occupied. Red arrows represent dirac deltas standing for an ideal delay-Doppler distribution (no BT limitation).
@@
Figure 17: Estimated delay-Doppler distribution, frequency-time aperture limited in BT, sparsely occupied resource grid.
@@
Figure 18: Sparse sampling in the BT aperture constrained $(f,t)$ OFDM resource grid. The rippled surface is $|H(f,t)|$ for the two-path propagation model in (8) and (9). The red rectangles indicate the active resource blocks where $H(t,f)$ is sampled.
@@
Figure 19: Sparse sampling of $H(t,f)$. The sampling pattern shows empty carriers (noise only) and carriers with different magnitudes because of multilevel modulation. The complex sinusoid (blue and green) depicts the response of a single path in the $(t,f)$ domain whose parameters frequency, phase and magnitude have to be estimated.
@@
The formulated signal model (8) also provides a theoretical basis for KPIs that quantify the accuracy and resolution of the parameter estimates. This allows, for example, to predict the estimation performance based on available REs, but also to judge what REs are most beneficial to solve the estimation problem. This is important for allocation of resource blocks in the OFDMA frame. The simplest way to assess the performance of an estimator is to quantify its bias, i.e., the systematic estimation error, and its variance, i.e., the statistical distribution of the estimates. Luckily, the often-employed maximum likelihood estimators for propagation parameter estimation [ref64, ref67] possess the beneficial statistical features of unbiasedness and efficiency, such that the assessment of estimation performance reduces to analyzing theoretical lower bounds for the estimation variance. That is why the Cramér-Rao Bound (CRB) is frequently employed as design criterion for waveform optimization judging the achievable estimation variance. The analytical expressions for this bound aid for example in the design of waveforms explicitly tailored to the OFDM-related estimation problem given here [gedschold2025waveformDesignVerification, ref36, Eijnde1991_SignalDesign]. The CRB results from the inverse of the so-called Fisher information matrix
@@
For conciseness, all model parameters are summarized in $\bm{\theta}=[\bm{\tau}_{p},\bm{\alpha}_{p},\bm{\gamma}_{p}]$. Conveniently, the Slepian-Bangs formula [stoica2005slepian] allows to directly calculate the Fisher information matrix from the partial derivatives of the of the signal model in the time-frequency aperture domain with respect to the target parameters in Doppler-Delay.
@@
Here, we consider $\bm{H}$ as a sampled, i.e., finite dimensional, version of $H(f,t)$ (see Figure 18) and it can be understood as discrete in the carrier and symbol dimensions, sampled by the sparse resource blocks as indicted in Figure 18. $\langle\cdot,\cdot\rangle$ is the inner product, $\mathcal{R}$ the real part, and $\sigma^{2}$ the variance of an additive white Gaussian noise process.
@@
There are different possibilities how to get knowledge about $\bm{H}$. In the most simple case of one target (one propagation path, no superposition of multiple paths) $H(f,t)$ is not frequency selective (constant magnitude). In this case the CRB can be calculated only by assuming a certain bandwidth and SNR. This can be used for allocation of resource blocks in the OFDMA frame as will be discussed below. For this end we can directly assess the contribution of each carrier or resource block to the final parameter estimate. Since the Fisher information is additive over the samples of the signal model [javorszky1996] (resulting from the inner product), we can make the same comparison also for several superimposed paths. However, now $H(f,t)$ becomes frequency selective and the Fisher information will be subjected to some variance. In a two cycle estimation procedure we can get an initial estimate of $H(f,t)$ from pilot based CSI estimates which enhances the estimation in the next step or we make predictive use of former estimates which continuously updates the predistortion along the observed track.
@@
The simple single target case leads to the result, that delay parameter estimation benefits from concentrating transmit power on the edges of the available spectrum [ref36, Mura2025waveformDesignOFDM_ISAC] as already intuitively stated in the section on MS Radio Access and Scheduling. The same is also known from radar target estimation, where the lower bound for the estimation variance is proportional to the inverse of the root mean squared (RMS)-bandwidth of the signal [richards2014RadarSignalProcessing, Ch. 7.2.1]. In contrast to the plain bandwidth definition, the RMS-bandwidth is the second-order moment of the power spectral density of the signal. Hence, the RMS-bandwidth increases when concentrating more transmit power on the edges of the signal’s spectrum.
@@
The key takeaways from this subsection are as follows. (i) Range-Doppler resolution of target parameters resembles a 2D harmonic retrieval problem, which extends to higher dimensions when spatial dimensions are added. (ii) Multidimensional estimation should be solved as a joint estimator rather than a sequential 1D estimator, as this avoids any parameter association problems. In addition, resolution in dimension with a small aperture can benefit from the resolution of another aperture dimension. Finally joint multi-dimensional resolution counts. (iii) Estimation must be able to handle the sparse distribution of resource elements in $(f,t)$ due to OFDMA and TDD. (iv) Model-based estimation is the preferred approach for high-resolution estimation of radar parameters from finite and sparsely observed apertures [ref64, ref67, miranda2025joint_est_ofdma, miranda2025mbpe_sparse_ofdma]. (v) Performance figures based on the CRB can be used to allocate and control the resource elements of multiple sensing links in the OFDM frame according to the aspired detection, resolution and tracking performance (see also subsection below).
@@
The fundamental advantage of model based estimation is its physical relevance. The model consists of two parts. The propagation data model in (9) reflects the structural information of multipath in terms of delay and Doppler. This can be seen as a priori information about the physical process to be identified. The unknown information is contained in the set of parameters to be estimated. The propagation data model can be extended to include propagation directions and refined, e.g., to include coupling between parameters such as delay and Doppler or DoA and delay, etc. [ref67]. The second part is the device data model that describes the measurement device and the process of data acquisition. In addition to the basic aperture parameters, bandwidth and observation time (that are considered in delay-Doppler estimation), it can include various linear and nonlinear static and time-variant ISAC sensor response functions (at Tx and Rx). This opens the possibility to consider any hardware impairments, far beyond just bandwidth and observation time constraints [ref35]. Also specific waveform features resulting from the primary communication function of ISAC can be taken into account. This topic will be continued in subsection Spatial Precoding, Beamforming, and Bidirectional Estimation below when directional filtering and estimation (DoA, DoD) is discussed.

### v2

@@ V-C Recovery of the Correlation Reference Signal
Another approach to filter out the remaining multipath is beamforming, which is discussed in the context of spatial precoding later on.
V-D Model-based Bistatic Target Parameter Estimation
@@
This subsection deals with the estimation of target-related parameters from observed ISAC OFDM frames. As described above, the 3D dynamic target state vector may be estimated from a set of concurrent path parameters in delay and Doppler. This is not a restriction on the generality, as we can easily add beamforming and even directional estimation, as will be discussed later on. The aim here is to estimate the statistical delay-Doppler distribution, which is also known as the 2D point spread function of linear time-variant systems. In Bello’s classical terminology [ref63], it is called Scattering Function (ScF). This is the magnitude squared of the Spreading Function (SF) $S(\tau,\alpha)$ that results from the inverse 2D Fourier transform of the time-variant transfer function $H(f,t)$.
@@
$\chi(\tau,\alpha)=\int\limits_{-\infty}^{+\infty}x(t)x^{\ast}(t-\tau)\exp(\jmath 2\pi\alpha t)\mathrm{d}t\,.$ (6)
The AF describes the resolution capability of the excitation waveform associated with its impulse compression performance and correlation gain. While the ideal AF according to (6) is a 2D Dirac delta function (”thumbtack”), for uniformly occupied OFDM frequency-time frames it appears as a 2D $\operatorname*{sinc}$ function because of its limited bandwidth $B$ and frame duration $T$. Although the side lobes of this AF are often considered as disturbing, the main lobe is the narrowest, and the correlation gain is the highest for a given area of support $BT$, in the frequency-time domain. The resulting detector may therefore be seen as optimum in terms of maximum likelihood performance in case one path is to be estimated, simple FFT processing, and constant noise energy distribution in the delay domain.
@@
The first equality is the classical Wiener-Hopf equation. It is also known as Wiener-deconvolution. In the general framework of frequency domain system identification [ref30], estimation of the auto / cross power spectral densities $R_{X}(f)$ and $R_{Y,X}(f)$ of stationary random signals would need a separate statistical averaging procedure. However, it follows directly from the OFDM paradigm that the orthogonality of the received OFDM carriers is kept, and any leakage noise of the estimated samples in the frequency domain is avoided. This results from the Cyclic Prefix (CPX), which is added before transmission and removed at the receiver before FFTs processing. The absence of further statistical averaging allows for rigorous simplification, which finally leads to the well-known inverse filter approach, which consists of dividing by the recovered correlation reference signal as summarized in equation (7). Actually, PCL applies a similar processing scheme [ref38]. This also corresponds to standard zero-forcing equalization. Another reason that allows for the factorization of the frequency / delay and time / Doppler transform is related to the “underspread” characteristics of the mobile radio channel [ref95]. This means it is a slowly time-variant system, which can be handled as block-wise time-invariant as long as the OFDM symbols are chosen to be short enough to avoid time-selective fading (the block fading assumption). This finally allows the factorization, which avoids full 2D FFTs and matches perfectly to the OFDM receiver processing as depicted in Figure 15. The adjacent Doppler analysis is carried out by a subsequent block of FFTs along slow time over $M$ stacked CIRs. These FFTs can also be interpreted as a Doppler filter bank. The number of transformed consecutive OFDM symbols (the OFDM frame) determines the coherent radar integration time $T$. Further noncoherent integration can be added.
@@
Figure 15: Basic ISAC-OFDM receiver signal processing in frequency-time domain and estimation of the scattering function. The blue marked area indicates standard OFDM processing.
Although the simple 2D FFT based frequency domain system identification scheme has several advantages, it also has some drawbacks and pitfalls. Among those are: (i) Even if the OFDM frames were completely filled with carriers of equal strength, the strong sinc sidelobes in the 2D delay/Doppler plane may mask weak reflections. The resolution of components of similar strength is limited by the AF mainlobe width, which is given by the $1/B\times 1/T$ aperture size, where $B$ is the bandwidth and $T$ is the coherent processing interval in slow time. (ii) The estimated SF is calculated at the predefined FFT grid in the $(\tau,\alpha)$ plane. However, the maxima of $\left|S(\tau,\alpha)\right|^{2}$, which mark the $(\tau,\alpha)$ target parameters, are not bound to this grid. Therefore, we would need efficient local interpolation schemes if we apply the FFT. (iii) Dynamic targets that are revealed by a non-zero Doppler shift will also change their observed delay (resp. range). Therefore, the estimated SF may be blurred by range-Doppler migration if the delay resolution (i.e., bandwidth) is high and the chosen coherent integration time is too long, since the Doppler filter bank assumes constant delays. (iv) The most fundamental drawback, however, comes into play when we are faced with a sparse, non-uniform distribution of the REs in the frequency / time frame that is so important for ISAC as discussed above. In this case, the AF is distorted in an uncontrolled manner, which may cause detection ambiguity and make correct estimation impossible. We should be aware that application of non-rectangular window functions (tapering) in frequency and / or slow time aperture space is always suboptimal since that reduces resolution and increases noise induced variance. They also do not help at all in case of a sparse RE distribution.
In the sequel, we describe an example that explains the problems of Fourier Transform (FT) based estimation of the ScF that arise with aperture-limited and sparse observations in the frequency / slow time domain. At the same time, we introduce a model-based estimation approach, which can solve both problems. We consider a propagation situation which is determined by $P$ specular paths with delay / Doppler parameters $\tau_{p}$, $\alpha_{p}$ and path weight $\gamma_{p}$, respectively
@@
where $H(f,t)$ denotes the slowly time-variant radio channel’s system response, limited to finite observation apertures, both in frequency and (slow) time. After a 2D FT, we obtain the complex SF $S(\tau,\alpha)$ which is affected by a $\operatorname*{sinc}$ function in both dimensions. The symbol $\ast$ is a shortcut for a convolution operation in the dimensions $\tau$ and $\alpha$, respectively.
@@
Here, $\left|S(\tau,\alpha)\right|^{2}$ is the so-called ScF, where the detection takes place. It is also referred to as the delay-Doppler distribution. For the example $P=2$, the red arrows in Figure 16 indicate the ideal case of two Diracs if no bandwidth-time (BT) limitation in the $(f,t)$ aperture plane takes place. The rippled surface is the magnitude of $H(f,t)$ according to (8) that results from the superposition of the two equal power propagation paths. The complex valued $H(f,t)$ is sampled and used for the FFT transform. The 2D rectangular limited support in the $(f,t)$ aperture domain describes the observed data taken by the OFDM frame. The effect of BT limitation is also included in (8) and (9). If the OFDM carriers are uniformly power loaded, this appears to be the optimal processing scheme. Some interpolation in delay-Doppler for better detection of maxima can be achieved by zero filling in the $(f,t)$ aperture space outside the observed OFDM frame. Although the two paths appear to be smeared in the delay-Doppler domain by the 2D convolution with a 2D $\operatorname*{sinc}$ function (see Figure 16 ), they can be easily identified since they are well separated.
@@
Now we consider a OFDMA resource grid with sparsely allocated REs as indicated by the red dots in the $(f,t)$-frame in Figure 18. $H(f,t)$ is now effectively sampled by the sparsely allocated REs. The missing samples due to empty carriers are set to zero for FFT-processing. This is common practice since the FFT requires a uniform sample grid. However, simply replacing samples that are not measured with zeros is a bad choice, as zeros are definitely wrong. The resulting estimated delay-Doppler distribution is therefore subject to corresponding distortions, which are characterized by uncontrolled side lobes and can make detection impossible as depicted by the example in Figure 17.
@@
Fortunately, the solution is obvious. A quick look at (8) explains that we can simply consider the equation given there for the analytical description of the problem as a parameterized model of the complex transfer function. So we just need to estimate the parameters of this model. If we have enough well-allocated REs in the $(f,t)$ resource frame, we can estimate the model parameters by fitting the $H(f,t)$ model at the measured RE positions. This is very well supported by the orthogonality of the system response in the frequency domain with respect to the OFDM carriers. Parameter estimation is achieved by minimizing a properly defined least squares cost function. In some general sense, this estimation procedure constitutes a multidimensional maximum likelihood parameter estimation scheme [ref64, ref67]. A deep learning based approach for joint delay-Doppler model parameter optimization from frequency and time samples can be found here [schieler2025prop_est_cnn]. Efficient real-time implementation is another problem and remains an active area of research. While batch processing may not be feasible, recursive estimation seems to be more appropriate. It is also worth mentioning that the model order $P$ is, in most cases, not known a priori and has to be determined. Finally, the complex path weights $\gamma_{p}$ are estimated. Moreover, as $\gamma_{p}$ is polarization sensitive, it can be given by a $2\times 2$ Jones matrix that accounts for HH, VV co-polar and HV, VH cross-polar I/O responses.
@@
Model-based estimation allows us to obtain the set of discrete delay-Doppler parameters ($\tau$, $\alpha$) without having to apply an FFT. This avoids any aperture truncation and sparsity gap errors, which would reduce resolution and lead to spurious sidelobes, as discussed above. Figure 18 explains that the propagation data model effectively interpolates between the observed sparse samples in $(f,t)$, thus avoiding the artifacts that would result from a simple FFT application. At the same time, the effect of high resolution is explained, since the model function implicitly extrapolates the observed data beyond the $BT$ limits, where $H(f,t)$ is measured, thus evading the Fourier resolution limit (beyond Rayleigh resolution). A key aspect of this estimation procedure is that it decomposes the relevant propagation components in the complex linear signal domain. This way, their mutual interaction in the delay-Doppler domain, because of limited support and sparsity in the observed frequency-time aperture domain, is removed. The cancellation of overlapping non-resolved contributions and sidelobes in the complex signal domain seems to be more powerful than the well-known CLEAN procedure [hoegbom1974clean], which works in the real-valued point spread or correlation domain.
@@
However, when considering the waveform of the data payload as an excitation signal for system identification, further considerations are necessary. Figure 19 shows samples of $H(f,t)$ at some cross section in $f$ or $t$. The problems are as follows: Empty REs have to be identified and excluded from the inverse filter operation. A null hypothesis test allows empty but noisy resource elements to be detected and removed from the estimation. The corresponding gap positions are effectively filled by interpolation. Another problem relates to the non-uniform power level of data carriers due to stronger CSI pilots and because of adaptive power loading and multi-level QAM coding. While the simple inverse filter would be affected by uneven noise levels, the SNR of the carriers can be estimated at the receiver and considered in parameter estimation as weighted least squares fitting, which gives the samples with lower SNR a smaller weight.
@@
Figure 16: Observed delay-Doppler distribution. The frequency-time resource grid is limited in BT and fully occupied. Red arrows represent Dirac deltas standing for an ideal delay-Doppler distribution (no BT limitation). Calculated from 48x48 carriers in frequency and time, padded with zeros for 512x512 FFT and interpolation in the delay-Doppler domain.
@@
Figure 17: Estimated delay-Doppler distribution, frequency-time aperture limited in BT, sparsely occupied resource grid. Calculated with the same FFT parameters as in Figure 16.
@@
Figure 18: Sparse sampling in the BT aperture-constrained $(f,t)$ OFDM resource grid. The rippled surface is $|H(f,t)|$ for the two-path propagation model in (8) and (9). The red rectangles indicate the active resource blocks where $H(f,t)$ is sampled.
@@
Figure 19: Sparse sampling of $H(f,t)$. The sampling pattern shows empty carriers (noise only) and carriers with different magnitudes because of multilevel modulation. The complex sinusoid (blue and green) depicts the response of a single path in the $(f,t)$ domain whose parameters frequency, phase, and magnitude have to be estimated.
@@
The formulated signal model (8) also provides a theoretical basis for KPIs that quantify the accuracy and resolution of the parameter estimates. This allows, for example, to predict the estimation performance based on available REs, but also to judge which REs are most beneficial to solve the estimation problem. This is important for allocation of resource blocks in the OFDMA frame. The simplest way to assess the performance of an estimator is to quantify its bias, i.e., the systematic estimation error, and its variance, i.e., the statistical distribution of the estimates. Luckily, the often-employed maximum likelihood estimators for propagation parameter estimation [ref64, ref67] possess the beneficial statistical features of unbiasedness and efficiency, such that the assessment of estimation performance reduces to analyzing theoretical lower bounds for the estimation variance. That is why the CRB is frequently employed as a design criterion for waveform optimization, judging the achievable estimation variance. The analytical expressions for this bound aid, for example, in the design of waveforms explicitly tailored to the OFDM-related estimation problem given here [gedschold2025waveformDesignVerification, ref36, Eijnde1991_SignalDesign]. The CRB results from the inverse of the so-called Fisher information matrix
@@
For conciseness, all model parameters are summarized in $\bm{\theta}=[\bm{\tau}_{p},\bm{\alpha}_{p},\bm{\gamma}_{p}]$. Conveniently, the Slepian-Bangs formula [stoica2005slepian] allows us to directly calculate the Fisher information matrix from the partial derivatives of the signal model in the time-frequency aperture domain with respect to the target parameters in delay-Doppler
@@
Here, we consider $\bm{H}$ as a sampled, i.e., finite-dimensional, version of $H(f,t)$ (see Figure 18), and it can be understood as discrete in the carrier and symbol dimensions, sampled by the sparse resource blocks as indicated in Figure 18. $\langle\cdot,\cdot\rangle$ is the inner product, $\mathcal{R}$ the real part, and $\sigma^{2}$ the variance of an additive white Gaussian noise process.
@@
There are different possibilities for getting knowledge about $\bm{H}$. In the simplest case of one target (one propagation path, no superposition of multiple paths), $H(f,t)$ is not frequency selective (constant magnitude). In this case, the CRB can be calculated only by assuming a certain bandwidth and SNR. This can be used for allocation of resource blocks in the OFDMA frame, as will be discussed below. For this end, we can directly assess the contribution of each carrier or resource block to the final parameter estimate. Since the Fisher information is additive over the samples of the signal model [javorszky1996] (resulting from the inner product), we can also make the same comparison for several superimposed paths. However, now $H(f,t)$ becomes frequency selective, and the Fisher information will be subjected to some variance. In a two-cycle estimation procedure, we can get an initial estimate of $H(f,t)$ from pilot-based CSI estimates, which enhances the estimation in the next step, or we make predictive use of former estimates, which continuously updates the predistortion along the observed track.
@@
The simple single target case leads to the result that delay parameter estimation benefits from concentrating transmit power on the edges of the available spectrum [ref36, Mura2025waveformDesignOFDM_ISAC] as already intuitively stated in the section on MS Radio Access and Scheduling. The same is also known from radar target estimation, where the lower bound for the estimation variance is proportional to the inverse of the root mean squared (RMS)-bandwidth of the signal [richards2014RadarSignalProcessing, Ch. 7.2.1]. In contrast to the plain bandwidth definition, the RMS-bandwidth is the second-order moment of the power spectral density of the signal. Hence, the RMS-bandwidth increases when concentrating more transmit power on the edges of the signal’s spectrum.
@@
The key takeaways from this subsection are as follows. (i) Range-Doppler resolution of target parameters resembles a 2D harmonic retrieval problem, which extends to higher dimensions when spatial dimensions are added. (ii) Multidimensional estimation should be solved as a joint estimator rather than a sequential 1D estimator, as this avoids any parameter association problems. In addition, resolution in a dimension with a small aperture can benefit from the resolution of another aperture dimension. Finally, joint multi-dimensional resolution counts. (iii) Estimation must be able to handle the sparse distribution of resource elements in $(f,t)$ due to OFDMA and TDD. (iv) Model-based estimation is the preferred approach for high-resolution estimation of radar parameters from finite and sparsely observed apertures [ref64, ref67, miranda2025joint_est_ofdma, miranda2025mbpe_sparse_ofdma]. (v) Performance figures based on the CRB can be used to allocate and control the resource elements of multiple sensing links in the OFDM frame according to the aspired detection, resolution, and tracking performance (see also subsection below).
@@
The fundamental advantage of model-based estimation is its physical relevance. The model consists of two parts. The propagation data model in (9) reflects the structural information of multipath in terms of delay and Doppler. This can be seen as a priori information about the physical process to be identified. The unknown information is contained in the set of parameters to be estimated. The propagation data model can be extended to include propagation directions and refined, e.g., to include coupling between parameters such as delay and Doppler or DoA and delay, etc. [ref67]. The second part is the device data model that describes the measurement device and the process of data acquisition. In addition to the basic aperture parameters, bandwidth, and observation time (which are considered in delay-Doppler estimation), it can include various linear and nonlinear static and time-variant ISAC sensor response functions (at Tx and Rx). This opens the possibility to consider any hardware impairments, far beyond just bandwidth and observation time constraints [ref35]. Also, specific waveform features resulting from the primary communication function of ISAC can be taken into account. This topic will be continued in subsection Spatial Precoding, Beamforming, and Bidirectional Estimation below when directional filtering and estimation (DoA, DoD) is discussed.